# Bridging Training and Execution via Dynamic Directed Graph-Based Communication in Cooperative Multi-Agent Systems

### v1

@@ 1 Introduction
Cooperative multi-agent reinforcement learning (MARL) problems have emerged in the past decade as a vital framework for addressing intricate collaborative challenges in real-world scenarios (Wei et al. 2019; Vinyals et al. 2019), attracting considerable attention and showcasing substantial potential for practical applications and commercial viability. A natural approach to cooperative MARL regards the multi-agent system as a whole using centralized methods (Tan 1993). A natural approach to cooperative MARL regards the multi-agent system as a whole using centralized methods(Tan 1993), which face scalability challenges and limitations associated with centralized controllers (Foerster et al. 2018). Another approach is decentralized, where each agent learns its policy with single agent techniques (Tampuu et al. 2017), addressing the limitations of centralized controllers but introducing non-stationarity and credit assignment problems (Lowe et al. 2017). To mitigate these issues, decentralized policies can be learned using the paradigm of centralized training with decentralized execution (CTDE) (Sunehag et al. 2017; Yuan et al. 2022; Yu et al. 2022; Zhang, Zhang, and Lin 2019), which includes value-based (Rashid et al. 2020; Eccles et al. 2019; Wang et al. 2021) and policy-based (Yu et al. 2022; Lin et al. 2021) frameworks. CTDE-based MARL algorithms rely on access to the global state, which is often an idealized assumption.
@@
In this work, we focus on developing communication mechanisms in MARL to enhance team performance on collaborative tasks while minimizing communication costs and avoiding reliance on global state information. We propose a transformer-based graph coarsening network (TGCNet), a novel communication-based MARL algorithm. In TGCNet, all cooperative agents are modeled as dynamic directed graphs, with each agent as a node. A transformer-based multi-key gated communication mechanism is integrated with the Q-network, allowing for end-to-end training without additional loss functions. This communication network learns the gating mechanism through a hard additive attention mechanism (Bahdanau, Cho, and Bengio 2014). By setting multi-key for aggregation, the number of communications between agents at each time step is learned, and an encoder fuses the communicated information. We use a transformer (Vaswani et al. 2017) architecture to process this fused information, from which weights are integrated and extracted. This approach generates an adjacency trajectory matrix that represents the communication connectivity structure between agents in the dynamic directed graphs. Meanwhile, we propose a novel communicative and cooperative MARL paradigm that connects training and execution through dynamic directed graphs, enabling simultaneous communication during training and execution phases. In addition, we theoretically decompose the global value function $Q_{\mathrm{tot}}$ into individual value functions $Q^{i}$, using each agent’s local observation and adjacency trajectory matrix as inputs. Building upon these theoretical results, we propose a practical graph coarsening network that employs self-attention pooling and coarsening operations to approximate the global state. We conduct a comprehensive empirical study across diverse cooperative multi-agent benchmarks to evaluate the efficacy of our proposed methodology. Specifically, we evaluate our method on the Hallway scenario (Wang et al. 2019), the Level-Based Foraging environment (Rangwala and Williams 2020), and eight distinct maps from the StarCraft Multi-Agent Challenge (SMAC) (Samvelyan et al. 2019). In addition, we perform component analyses to demonstrate the effectiveness of the multi-key gated communication network and the graph coarsening network. The contributions of this study are as follows:
  • We formalize a novel general paradigm for communicative, cooperative MARL, which bridges training and execution through dynamic directed graphs. The communication policy is represented by the structure of dynamic directed graphs, which is shared during both training and execution.
@@ 2 Related Works
Communication is vital for MARL to capture the dependencies between agent actions. In addition, it has been demonstrated to effectively improve exploration and team rewards (Apicella et al. 2012). Our work builds on and relates to previous research on MARL and communication mechanisms. The existing MARL can be divided into two categories based on whether or not a communication mechanism is set up. MARL without communication uses the CTDE paradigm, which has been successfully implemented with value-based and policy-based algorithms. The focus of policy-based MARL is to stabilize training with centralized state value estimation. Representative works in this area include COMA (Foerster et al. 2018), MADDPG (Lowe et al. 2017), and MAPPO (Yu et al. 2022). Another category is value-based MARL, following the Individual-Global-Max (IGM) principle; it focuses on value function factorization. This approach includes VDN (Sunehag et al. 2017), QMIX (Rashid et al. 2020), QTRAN (Son et al. 2019), and QPlex (Wang et al. 2020). Although these algorithms have shown significant performance in many multi-agent cooperative tasks, their effectiveness relies heavily on the introduction of global state and the setup of centralized trainer. Unlike existing works, our algorithm facilitates communication among agents, thereby eliminating reliance on global state.
@@
The second category focuses on the use of communication mechanisms to compensate for the limitations of local observations. Previous works enforced fixed and static broadcasting for communication. RIAL and DIAL (Foerster et al. 2016) were applied to improve communication skills by broadcasting messages across time steps. CommNet (Sukhbaatar, Fergus et al. 2016) broadcasted hidden states as messages and obtained fused information by averaging. IC3Net (Singh, Jain, and Sukhbaatar 2018) included a gating mechanism for learning when to broadcast using CommNet. BiCNet (Peng et al. 2017) and ATOC (Jiang and Lu 2018) considered bidirectional communication and used bidirectional RNNs (Zaremba, Sutskever, and Vinyals 2014) or LSTMs (Greff et al. 2016) to achieve bidirectional broadcast. Recent studies aim to develop more sophisticated communication mechanisms to prevent indiscriminate message broadcasting. DGN (Jiang et al. 2018) leveraged graph convolutional networks with relational kernels to capture dynamic agent interactions. TarMAC (Das et al. 2019) adopted an attention mechanism to determine whether two agents must communicate and distinguished the importance of incoming messages. G2ANet (Liu et al. 2020) constructed a communication interaction graph using a two-stage attention mechanism. However, these algorithms only facilitate communication before agents take actions, neglecting communication during the centralized training process. We propose a new multi-key gated communication network for bidirectional, modeling communication as a dynamic directed graph, combined with transformer for information fusion and retrieval, to address these concerns. Moreover, we integrate graph neural networks to convert the communication flow into a structured network of dynamic directed graphs that change dynamically at each time step, and leverage graph coarsening network to approximate the global state.
Figure 1: (a) Adjacency trajectory matrix at time $t$. (b) Dynamic directed graph. The dynamic directed graph can represent the associations and communication structures between nodes at a certain moment.
@@ 3 Preliminaries
We model cooperative multi-agent systems as Dec-POMDP (Oliehoek and Amato 2016), which imposes partially observable settings on multi-agent Markov decision processes. A Dec-POMDP can be described by a tuple $\langle\mathcal{N},\mathcal{S},\mathcal{A},P,\Omega,O,R,\gamma,\mathcal{M}\rangle$, where $\mathcal{N}=\{1,\cdots,n\}$ indicates the set of $n$ agents, $\mathcal{S}$ denotes the set of global states, $\mathcal{A}$ refers to the set of actions, $P$ represents the state transition function, $\Omega$ denotes the set of observations, $O$ refers to the observation function, $R$ represents the reward function, $\gamma\in[0,1)$ indicates the discount factor, and $\mathcal{M}$ is the set of messages that agents can communicate. At each time step $t$, each agent $i\in\mathcal{N}$ receives an observation $o_{t}^{i}\in\Omega$ from the observation function $O(s_{t},i)$, where $s_{t}\in\mathcal{S}$. Each agent follows an individual policy $\pi^{i}(a_{t}^{i}|\tau_{t}^{i},m_{t}^{i})$, where $\tau_{t}^{i}=(o_{1}^{i},a_{1}^{i},\ldots,o_{t-1}^{i},a_{t-1}^{i},o_{t}^{i})$ is the action-observation history of agent $i$, and $m_{t}^{i}\in\mathcal{M}$ is the message received by agent $i$ at time $t$. The joint action $\boldsymbol{a}_{t}=\langle a_{t}^{1},\ldots,a_{t}^{n}\rangle$ leads to the next state $s_{t+1}\sim P(s_{t+1}|s_{t},\boldsymbol{a}_{t})$, and the team receives a global reward $R(s_{t},\boldsymbol{a}_{t})$. The objective is to find a joint policy $\boldsymbol{\pi}=\langle\pi^{1},\ldots,\pi^{n}\rangle$ that maximizes the global value function $Q_{\mathrm{tot}}^{\boldsymbol{\pi}}(\boldsymbol{\tau},\boldsymbol{a})=\mathbb{E}_{s,\boldsymbol{a}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s,\boldsymbol{a})\mid s_{0}=s,\boldsymbol{a}_{0}=\boldsymbol{a},\boldsymbol{\pi}\right]$, with $\boldsymbol{\tau}=\langle\tau_{1},\ldots,\tau_{n}\rangle$.
@@
Based on Dec-POMDP, we propose a communication topology using dynamic directed graphs. The structural properties of the graph are represented by the adjacency trajectory matrix $\mathbf{A}$, which is denoted as $\mathbf{A}\in\mathbb{B}^{l\times n\times n}$, where $\mathbb{B}$ indicates the Boolean matrix, $l$ denotes the length of the trajectory and $n$ indicates the number of agents. If $\lceil A_{t}^{ij}\rceil=0$, then no communication occurred from agent $i$ to agent $j$ at time $t$. If $\lceil A_{t}^{ij}\rceil=1$, then communication transpired from agent $i$ to agent $j$ at time $t$. As an example, the adjacency trajectory matrix $A_{t}$ at time $t$ is shown in Figure 1(a), and the corresponding dynamic directed graph is shown in Figure 1(b). After traversing the dynamic directed graph to agent $i$, the received message is computed as $\tilde{m}_{t}^{i}=A_{t}^{i}\odot m_{t}^{i}$, where $\odot$ means element-wise multiplication. For value-based MARL, which uses the action-value function to update policies, where in the distributed execution phase, each agent learns a Q-Network $Q^{i}(\tau,a,\tilde{m};\theta^{i})$ (Mnih et al. 2015) to approximate the action-value function $Q^{i}(s,a)$. In the centralized training phase, the group of agents learns a mix network $Q_{\mathrm{tot}}(\boldsymbol{\tau},\boldsymbol{a},\boldsymbol{\tilde{m}},s;\boldsymbol{\theta})$ to approximate the global action value function $Q_{\mathrm{tot}}(s,\boldsymbol{a})$. The parameters $\boldsymbol{\theta}$ are learned by minimizing the expected temporal difference (TD) error:
$\mathcal{L}(\theta)=\sum_{i=1}^{b}\left[\left(y_{i}^{\mathrm{tot}}-Q_{\mathrm{tot}}(\boldsymbol{\tau}_{t},\boldsymbol{a}_{t},\boldsymbol{\tilde{m}}_{t},s_{t};\boldsymbol{\theta})\right)^{2}\right],$ (1)
@@ 4 Method
In this section, we introduce the detailed structure and design of TGCNet. Fundamentally, TGCNet is an extension of value-based MARL. Our original intention was for TGCNet to simultaneously solve the five communication problems for MARL, specifically in Section 1. Our motivation stems from the objective of communication in MARL, which is to mitigate the negative effects of limited observations in the Dec-POMDP. Through communication, agents can obtain an approximation of the global state. Our aim is to learn a state function $S(o^{i}_{t},m^{i}_{t},A^{i}_{t})$. Moreover, if an agent’s information significantly affects the global state for agent $i$ during centralized training, it should also be crucial during decentralized execution. Communication plays two roles in TGCNet. In the Q-network, the agent uses communication to acquire information beyond its immediate observations to assist in decision making, and the message $m_{t}^{i}$ at time $t$ defined as the action-observation history of all other agents except agent $i$, denoted as $\tau^{-i}_{t}$. In the mix network, agents coarsen and aggregate individual local observations into global states through communication and the message $m_{t}^{i}$ at time $t$ defined as the local observations of all other agents except agent $i$, denoted as $o^{-i}_{t}$. In addition, the network is scalable to scenarios where agents may change and can be plugged into any CTDE structure.
Figure 2: (a) Graph coarsening network and mix network. The inputs for the graph coarsening network include the local observations of the agents and the adjacency trajectory matrix. (b) Overall TGCNet architecture. (c) Transformer-Based multi-key gated communication mechanism. The communication mechanism outputs not only the individual state value function $Q^{i}(\tau^{i},a^{i})$ and hidden variables $h_{t}^{i}$ but also the adjacency trajectory matrix $A^{i}_{t}$. Then, this matrix is passed into the graph coarsening network to complete end-to-end backward propagation updates.
@@ 4.1 Theoretical Analysis
We first theoretically analyze how the state function is constructed in the mix network. We construct a 3D dynamic directed which is defined as $G=(T,V,E)$, where $t\in T$ is the set of the trajectory in the time dimension, $v^{i}\in V$ denotes the set of vertex, which is composed of agents and $e^{i,j}\in E$ denotes the set of edges connecting agent $i$ and agent $j$. A dynamic directed graph can be regarded as a composition of $T$ subgraphs $G_{t}=(V_{t},E_{t})$ organized along a temporal trajectory. We define the adjacency trajectory matrix to describe the connections (communication) associations between graph vertices. The adjacency trajectory matrices for the dynamic directed graph and subgraphs are denoted by $\mathbf{A}$ and $A_{t}$. When a node can access the observations of other nodes, it can approximate the global state. However, this information may be either independent or partially redundant. We define the relationship between the global state and local observation using the following equation:
@@
where $\text{Agg}\left(\cdot\right)$ indicates an aggregator function, $N(\nu^{i})$ represents the set of neighbor nodes of node $i$, $\text{Overlap}\left(\cdot\right)$ denotes an overlap function and $\parallel$ represents the operator of parallel concatenation, which connects two vectors in parallel. It is evident that $o_{t}^{j},\forall\nu^{j}\in N(\nu^{i})\parallel o_{t}^{i}$ is essentially equivalent to $\tilde{m}^{i}_{t}$. When dealing with aggregation and overlap functions, it is crucial to maintain consistent output results regardless of the order of neighboring nodes. We have designed the summation aggregation functions $\text{Agg}^{\text{sum}}$ and pooling overlap functions $\text{Overlap}^{\text{pool}}$, as follows:

### v2

@@ 1 Introduction
Cooperative multi-agent reinforcement learning (MARL) problems have emerged in the past decade as a vital framework for addressing intricate collaborative challenges in real-world scenarios (Wei et al. 2019; Vinyals et al. 2019; Cheng et al. 2024; Li et al. 2023), attracting considerable attention and showcasing substantial potential for practical applications and commercial viability. A natural approach to cooperative MARL regards the multi-agent system as a whole using centralized methods (Tan 1993), which face scalability challenges and limitations associated with centralized controllers (Foerster et al. 2018). Another approach is decentralized, where each agent learns its policy with single agent techniques (Tampuu et al. 2017), addressing the limitations of centralized controllers but introducing non-stationarity and credit assignment problems (Lowe et al. 2017). To mitigate these issues, decentralized policies can be learned using the paradigm of centralized training with decentralized execution (CTDE) (Sunehag et al. 2017; Yuan et al. 2022; Yu et al. 2022; Zhang, Zhang, and Lin 2019), which includes value-based (Rashid et al. 2020; Eccles et al. 2019; Wang et al. 2021) and policy-based (Yu et al. 2022; Lin et al. 2021) frameworks. CTDE-based MARL algorithms rely on access to the global state, which is often an idealized assumption.
@@
In this work, we focus on developing communication mechanisms in MARL to enhance team performance on collaborative tasks, while minimizing communication costs and avoiding reliance on global state information. We propose a novel MARL algorithm, the Transformer-based graph coarsening network (TGCNet), which models cooperative agents’ communication as dynamic directed graphs, with each agent as a node. The adjacency trajectory matrix represents the structure of the dynamic directed graph. We connect the training and execution phases through dynamic directed graphs, allowing agents to communicate during both stages. TGCNet integrates a Transformer-based (Vaswani et al. 2017) multi-key gated communication mechanism with the Q-network, enabling end-to-end training without the need for additional loss functions. The multi-key gated mechanism learns through a hard additive attention approach (Bahdanau, Cho, and Bengio 2014), using multi-keys for aggregation to determine the number of communications between agents. Based on the dynamic directed graph, we introduce a graph coarsening network that utilizes self-attention pooling and coarsening operations to approximate the global state. We conduct a comprehensive empirical study across diverse cooperative multi-agent benchmarks to evaluate the efficacy of our proposed methodology. Specifically, we evaluate our method on the Hallway scenario (Wang et al. 2019), the Level-Based Foraging (LBF) environment (Rangwala and Williams 2020), and eight distinct maps from the StarCraft Multi-Agent Challenge (SMAC) (Samvelyan et al. 2019). In addition, we perform component analyses to demonstrate the effectiveness of the multi-key gated communication network and the graph coarsening network. The contributions of this study are as follows:
  • We formalize a novel general paradigm for communicative and cooperative MARL, which bridges training and execution through dynamic directed graphs. The communication policy is represented by the structure of dynamic directed graphs, which is shared during both training and execution.
@@ 2 Related Works
Communication is vital for MARL to capture the dependencies between agent actions. In addition, it has been demonstrated to effectively improve exploration and team rewards (Apicella et al. 2012). Our work builds on and relates to previous research on MARL and communication mechanisms. The existing MARL can be divided into two categories based on whether or not a communication mechanism is set up. MARL without communication uses the CTDE paradigm, which has been successfully implemented with value-based and policy-based algorithms. The focus of policy-based MARL is to stabilize training with centralized state value estimation. Representative works in this area include COMA (Foerster et al. 2018), MADDPG (Lowe et al. 2017), and MAPPO (Yu et al. 2022). The other category is value-based MARL, following the Individual-Global-Max (IGM) principle; it focuses on value function factorization. This approach includes VDN (Sunehag et al. 2017), QMIX (Rashid et al. 2020), QTRAN (Son et al. 2019), and QPlex (Wang et al. 2020). Although these algorithms have shown significant performance in many multi-agent cooperative tasks, their effectiveness relies heavily on the introduction of global state and the setup of centralized trainer. Unlike existing works, our algorithm facilitates communication among agents, thereby eliminating reliance on global state.
@@
The second category focuses on the use of communication mechanisms to compensate for the limitations of local observations. Previous works enforce fixed and static broadcasting for communication. RIAL and DIAL (Foerster et al. 2016) are applied to improve communication skills by broadcasting messages across time steps. CommNet (Sukhbaatar, Fergus et al. 2016) broadcasts hidden states as messages and obtains fused information by averaging. IC3Net (Singh, Jain, and Sukhbaatar 2018) includes a gating mechanism to learn when to broadcast using CommNet. BiCNet (Peng et al. 2017) and ATOC (Jiang and Lu 2018) consider bidirectional communication and use bidirectional RNNs (Zaremba, Sutskever, and Vinyals 2014) or LSTMs (Greff et al. 2016) to achieve bidirectional broadcast. Recent studies aim to develop more sophisticated communication mechanisms to prevent indiscriminate message broadcasting. DGN (Jiang et al. 2018) leverages graph convolutional networks with relational kernels to capture dynamic agent interactions. TarMAC (Das et al. 2019) adopts an attention mechanism to determine whether two agents must communicate and distinguishes the importance of incoming messages. G2ANet (Liu et al. 2020) constructs a communication interaction graph using a two-stage attention mechanism. However, these algorithms only facilitate communication before agents take actions, neglecting communication during the centralized training process. We propose a new multi-key gated communication network for bidirectional, modeling communication as a dynamic directed graph, combined with Transformer for information fusion and retrieval, to address these concerns. Moreover, we integrate graph neural networks to convert the communication flow into a structured network of dynamic directed graphs that change dynamically at each time step, and leverage graph coarsening network to approximate the global state.
@@ 3 Preliminaries
We model cooperative multi-agent systems as Dec-POMDP (Oliehoek and Amato 2016), which imposes partially observable settings on multi-agent Markov decision processes. A Dec-POMDP can be described by a tuple $\langle\mathcal{N},\mathcal{S},\mathcal{A},P,\Omega,O,R,\gamma,\mathcal{M}\rangle$, where $\mathcal{N}=\{1,\cdots,n\}$ indicates the set of $n$ agents, $\mathcal{S}$ denotes the set of global states, $\mathcal{A}$ refers to the set of actions, $P$ represents the state transition function, $\Omega$ denotes the set of observations, $O$ refers to the observation function, $R$ represents the reward function, $\gamma\in[0,1)$ indicates the discount factor, and $\mathcal{M}$ is the set of messages that agents can communicate. At each time step $t$, each agent $i\in\mathcal{N}$ receives an observation $o_{t}^{i}\in\Omega$ from the observation function $O(s_{t},i)$, where $s_{t}\in\mathcal{S}$. Each agent follows an individual policy $\pi^{i}(a_{t}^{i}|\tau_{t}^{i},m_{t}^{i})$, where $\tau_{t}^{i}=(o_{1}^{i},a_{1}^{i},\ldots,o_{t-1}^{i},a_{t-1}^{i},o_{t}^{i})$ is the action-observation history of agent $i$, and $m_{t}^{i}\in\mathcal{M}$ is the message received by agent $i$ at time $t$. The joint action $\boldsymbol{a}_{t}=\langle a_{t}^{1},\ldots,a_{t}^{n}\rangle$ leads to the next state $s_{t+1}\sim P(s_{t+1}|s_{t},\boldsymbol{a}_{t})$, and the team receives a global reward $R(s_{t},\boldsymbol{a}_{t})$. The objective is to find a joint policy $\boldsymbol{\pi}=\langle\pi^{1},\ldots,\pi^{n}\rangle$ that maximizes the global action-value function $Q_{\mathrm{tot}}^{\boldsymbol{\pi}}(\boldsymbol{\tau},\boldsymbol{a})=\mathbb{E}_{s,\boldsymbol{a}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s,\boldsymbol{a})\mid s_{0}=s,\boldsymbol{a}_{0}=\boldsymbol{a},\boldsymbol{\pi}\right]$, with $\boldsymbol{\tau}=\langle\tau_{1},\ldots,\tau_{n}\rangle$. Each agent serves as both sender and receiver.
@@
Figure 1: Dynamic directed graph. (a) Adjacency trajectory matrix at time $t$. (b) Dynamic directed graph. The dynamic directed graph can represent the associations and communication structures between nodes at a certain moment.
Based on Dec-POMDP, we propose a communication topology using dynamic directed graphs. The structural properties of the graph are represented by the adjacency trajectory matrix $\mathbf{A}$, which is denoted as $\mathbf{A}\in\mathbb{B}^{l\times n\times n}$, where $\mathbb{B}$ indicates the Boolean matrix, $l$ denotes the length of the trajectory and $n$ indicates the number of agents. If $\lceil A_{t}^{ij}\rceil=0$, then no communication occurred from agent $i$ to agent $j$ at time $t$. If $\lceil A_{t}^{ij}\rceil=1$, then communication transpired from agent $i$ to agent $j$ at time $t$. As an example, the adjacency trajectory matrix $A_{t}$ at time $t$ is shown in Figure 1(a), and the corresponding dynamic directed graph is shown in Figure 1(b). After traversing the dynamic directed graph to agent $i$, the received message is computed as $\tilde{m}_{t}^{i}=A_{t}^{i}\odot m_{t}^{i}$, where $\odot$ means element-wise multiplication. For value-based MARL, which uses the action-value function to update policies, where in the distributed execution phase, each agent learns a Q-Network $Q^{i}(\tau,a,\tilde{m};\theta)$ (Mnih et al. 2015) to approximate the action-value function $Q^{i}(s,a)$. In the centralized training phase, the group of agents learns a mix network $Q_{\mathrm{tot}}(\boldsymbol{\tau},\boldsymbol{a},\boldsymbol{\tilde{m}},s;\boldsymbol{\theta})$ to approximate the global action value function $Q_{\mathrm{tot}}(s,\boldsymbol{a})$. The parameters $\boldsymbol{\theta}$ are learned by minimizing the expected temporal difference (TD) error:
$\mathcal{L}(\boldsymbol{\theta})=\sum_{i=1}^{b}\left[\left(y_{i}^{\mathrm{tot}}-Q_{\mathrm{tot}}(\boldsymbol{\tau}_{t},\boldsymbol{a}_{t},\boldsymbol{\tilde{m}}_{t},s_{t};\boldsymbol{\theta})\right)^{2}\right],$ (1)
@@ 4 Method
In this section, we introduce the detailed structure and design of TGCNet. Fundamentally, TGCNet is an extension of value-based MARL. Our original intention is for TGCNet to simultaneously solve the five communication problems for MARL, specifically in Section 1. Our motivation stems from the objective of communication in MARL, which is to mitigate the negative effects of limited observations in the Dec-POMDP. Through communication, agents can obtain an approximation of the global state. Our aim is to learn a state function $S(o^{i}_{t},m^{i}_{t},A^{i}_{t})$. Moreover, if an agent’s information significantly affects the global state for agent $i$ during centralized training, it should also be crucial during decentralized execution. Communication plays two roles in TGCNet. In the Q-network, the agent uses communication to acquire information beyond its immediate observations to assist in decision making, and the message $m_{t}^{i}$ at time $t$ defined as the action-observation history of all other agents except agent $i$, denoted as $\tau^{-i}_{t}$. In the mix network, agents coarsen and aggregate individual local observations into global states through communication and the message $m_{t}^{i}$ at time $t$ defined as the local observations of all other agents except agent $i$, denoted as $o^{-i}_{t}$. In addition, the network is scalable to scenarios where agents may change and can be plugged into any CTDE structure.
Figure 2: The network structure diagram of TGCNet. (a) Graph coarsening network and mix network. The inputs for the graph coarsening network include the local observations of the agents and the adjacency trajectory matrix. (b) Overall TGCNet architecture. (c) Transformer-Based multi-key gated communication mechanism. The communication mechanism outputs not only the individual state value function $Q^{i}(\tau^{i},a^{i})$ and hidden variables $h_{t}^{i}$ but also the adjacency trajectory matrix $A^{i}_{t}$. Then, this matrix is passed into the graph coarsening network to complete end-to-end backward propagation updates.
@@ 4.1 Theoretical Analysis
We first theoretically analyze how the state function is constructed in the mix network. We construct a 3D dynamic directed which is defined as $\mathbf{G}=(T,V,E)$, where $t\in T$ is the set of the trajectory in the time dimension, $v^{i}\in V$ denotes the set of vertex, which is composed of agents and $e^{i,j}\in E$ denotes the set of edges connecting agent $i$ and agent $j$. A dynamic directed graph can be regarded as a composition of $T$ subgraphs $G_{t}=(V_{t},E_{t})$ organized along a temporal trajectory. We define the adjacency trajectory matrix to describe the connections (communication) associations between graph vertices. The adjacency trajectory matrices for the dynamic directed graph and subgraphs are denoted by $\mathbf{A}$ and $A_{t}$. When a node can access the observations of other nodes, it can approximate the global state. However, this information may be either independent or partially redundant. We define the relationship between the global state and local observation using the following equation:
@@
where $\text{Agg}\left(\cdot\right)$ indicates an aggregator function, $N(\nu^{i})$ represents the set of neighbor nodes of node $i$, $\text{Overlap}\left(\cdot\right)$ denotes an overlap function and $\parallel$ represents the operator of parallel concatenation, which connects two vectors in parallel. It is evident that $o_{t}^{j},\forall\nu^{j}\in N(\nu^{i})\parallel o_{t}^{i}$ is essentially equivalent to $\tilde{m}^{i}_{t}$. When dealing with aggregation and overlap functions, it is crucial to maintain consistent output results regardless of the order of neighboring nodes. We design the summation aggregation functions $\text{Agg}^{\text{sum}}$ and pooling overlap functions $\text{Overlap}^{\text{pool}}$, as follows: